A superconducting solenoid is to be designed to have an interior field of 4.252 T with a maximum current of 1000 A. The permeability of free space is 1.25664 × 10−6 T · m/A. How many windings are required on a 1- meter solenoid length?
Answer in units of turns/m.
1
Expert's answer
2020-08-10T19:59:04-0400

As per the given question,

Magnetic field "(B)=4.252 T"

Maximum current "(i)=1000A"

"\mu_o =1.25664 × 10^{−6} T · m/A."

Interior magnetic field of the solenoid "(B)=\mu_o nI"

Number of binding per meter "(N)=\frac{B}{\mu_o I}"

Now, substituting the values in the above,

"N=\frac{4.25 T}{1.25664 × 10^{−6}\times 1000A.T · m/A.}"

"\Rightarrow N=3400 turn/meter"